# It Is Time To Steer: A Scalable Framework for Analysis-driven Attack Graph Generation

### v1

@@ Abstract
In modern computer networks where sophisticated cyber attacks occur daily, a timely cyber risk assessment becomes paramount. Attack Graph (AG) represents the best-suited solution to model and analyze multi-step attacks on computer networks, although they suffer from poor scalability due to their combinatorial complexity.
This paper introduces an analysis-driven framework for AG generation. It enables real-time attack path analysis before the completion of the AG generation with a quantifiable statistical significance. We further accelerate the AG generation by steering it with the analysis query and supporting a novel workflow in which the analyst can query the system anytime. To show the capabilities of the proposed framework, we perform an extensive quantitative validation and we present a realistic case study on networks of unprecedented size. It demonstrates the advantages of our approach in terms of scalability and fitting to common attack path analyses.
Index Terms:
Attack Graph, Attack Path Analysis, Progressive Computation, Statistical Significance, Computational Steering.
@@ 1 Introduction
In today’s digital world, where attackers are evolving smarter and the number of different vulnerabilities is growing, cyber-attacks are becoming more complex to mitigate and with a higher impact on organizations’ infrastructure [1]. In this scenario, it is crucial for any organization to promptly assess the potential cyber risks their networks are exposed to. This involves identifying the vulnerabilities in the network’s systems and quantifying the risk of their potential exploitation, which generally refers to the impact of one or more exploits and their likelihood to happen [2, 3, 4].
Given the growing complexity of attacks, as multi-step attacks [5] (i.e., sequential exploits of different vulnerabilities among network hosts), it is essential to model the threats in a network. Among the possible threat models [6, 7], Attack Graph (AG) [8, 9] is a graph-based representation of the potential attack paths in a network, used to analyze its cyber risk. For example, an analyst may be interested in determining the most critical hosts based on the number of paths traversing them. These kinds of analyses are named attack path analyses as they involve features of the attack paths [10].
Although AGs are very expressive attack models for attack path analysis, the first problem they show is the poorly scalable computation of the attack paths, even for networks of moderate size as AGs may grow exponentially in the size of network hosts and vulnerabilities [9]. Currently, providing scalable approaches for attack path analysis is still an open problem, and this hinders the possibility of promptly providing security strategies to harden the network. Different works addressed the problem of scalability of AG generation [11, 12, 13] and attack path analysis [14, 15, 16].
A second emerging problem is that they postulate a classic analysis workflow in which the generation process step comes first and the analysis comes next, aggravating the poor scalability of the generation process as it slows down or completely stops the analysis phase in realistic scenarios.
A final problem concerns the alignment between the current network situation and the related AG model: while the classic workflow permits the attack path analysis on the complete set of attack steps, it also implies that any change in the network needs a re-computation of the whole attack paths set to reflect such changes, exacerbating the scalability problem. These problems make it challenging to use AGs for today’s dynamic and large-scale networks.
@@
To overcome them and provide affordable analysis, we contribute a novel workflow for AG generation and attack path analysis. It leverages the concept of statistical significance [17] to express the degree of trustfulness of a partial AG, where not all the attack steps are computed. According to the foundational aspects of progressive data analysis [18, 19], it progressively supports attack path analysis during the AG generation, allowing preventive analyses with quantitative quality-controlled statistically significant results. To improve the accuracy of analysis performed on partial AGs, we design a steering approach which automatically accelerates the AG generation based on the requested attack path analysis. Given a path analysis query from an analyst, it models the queried attack path features, extracts from them the features of each single attack step, and uses them to prioritize the generation of attack paths that answer the analysis query. We evaluate the speed-up of this process through a comprehensive scalability analysis performed on hundreds of experiments, varying both the network size and vulnerability inventories. Finally, we present a realistic case study that shows the capabilities of the approach in networks whose size cannot be managed by state-of-the-art solutions.
Summarizing, this paper contributes: i) a novel framework for AG generation and attack path analysis, based on progressive data analysis; ii) a way to express statistical significance over partial AGs; iii) a novel approach to steer the AG generation based on attack path analysis queries; iv) a comprehensive validation through experimental evaluations for both statistically significant AG generation and the steering mechanism; v) a realistic case study demonstrating the possibility of managing large networks and conducting common attack path analyses.
@@ 2 Preliminaries
In this section, we present the fundamental notions of attack graphs and attack path analysis. In particular, we define their components and introduce the risk model used for the attack path analysis.
2.1 Attack Graph Model Definition
An Attack Graph (AG) is a graph-based representation of the possible paths an attacker can exploit to intrude and compromise an ICT network or, more generally, an ICT system. To generate it, two inputs are necessary [20, 21, 22]: (i) the network reachability graph (or network reachability matrix), and (ii) the network vulnerability inventory. A network reachability graph is a directed graph $R=(H,E)$ where nodes are associated with network hosts $h_{1},\cdots,h_{n}\in H$ and edges represent reachability conditions between hosts (taking into account firewall and routing rules). Thus, an edge $e(h_{1},h_{2})\in E$ indicates that host $h_{1}$ can communicate with host $h_{2}$ through a direct link.
A vulnerability inventory reports the list of network hosts with an associated set of vulnerabilities, typically obtained by combining the vulnerability knowledge bases, as the National Vulnerability Database (NVD) [23] by NIST [24], with vulnerability scanners (e.g., Nessus [25]).
With these inputs, AG modeling consists of the representation of all the dependencies between hosts and vulnerabilities that can be exploited by an attacker. Thus, the first step to model AGs is defining the semantics of attack nodes and edges to represent the structure of the attack steps. In this paper, we leverage the Topology Vulnerability Analysis (TVA) attack graph model [26], where the nodes are security conditions and represent the attacker access privileges in a specific host, while the edges are exploit dependencies and represent the possible movement of the attacker in case of successful vulnerability exploit. More formally:
@@ 2.1 Attack Graph Model Definition
An example of an AG is reported in Appendix B. Once the AG model is defined, the next step is the Attack Graph Generation, which is the computation of the attack paths to analyze the cyber risks of the potential attacks. Attack paths represent sequences of compromised devices and vulnerabilities exploited during the attack, and are used to compute scores or metrics like the likelihood of success of an attack, its impact, or the difficulty of execution [12, 14, 27].
@@
When generation ends, the resulting attack paths are used to estimate the risks of the different multi-step attacks, each one corresponding to a different path in the AG. The attack path risks may be used either to evaluate the overall exposure of the network (e.g., by aggregating the risks of the different paths to have a comprehensive risk evaluation) or to perform analyses according to specific requirements (e.g., analyzing which are the attacks that may cause more damages to the network).
Figure 1: Classic attack graph generation and analysis process and the progressive one. Dashed lines indicate periods in which the analysis is stalled. PAG=Partial Attack Graph, SPAG=statistically Significant Partial Attack Graph generation, PAP=Partial Attack Path analysis
2.2 Risk Model Definition
To evaluate the cyber risks of an attack path, we leverage the risk methodology designed by Gonzalez-Granadillo et al. [2]. They consider CVSS standard metrics [28] to estimate the likelihood and impact of an attack path $AP$, and calculate the risk according to its standard definition, that is the dot product of likelihood and impact [29].
$risk(AP)=likelihood(AP)\cdot impact(AP)$ (1)
In particular, the likelihood is calculated using the CVSS-3.1 exploitability metrics, which are Attack Vector (AV), Attack Complexity (AC), Privilege Required (PR), and User Interaction (UI). The impact is determined by the last node in the path, as it represents the attacker’s final goal, and it considers the CVSS-3.1 impact metrics, that are Confidentiality (C), Integrity (I), and Availability (A). Likelihood, impact, and risk are all in the range [0,1]. More details about the risk model are available in Appendix C.
According to the defined risk model, the risks of the different attack paths inform the analysis process and consequent decisions on the mitigation strategies. For this reason, we evaluate AG expressiveness through the distribution of the likelihood and impact of its paths.
@@ 2.2 Risk Model Definition
Let $AP$ be an attack path and let $u_{1,2},\cdots\in U_{AP}$ the sequence of its vulnerabilities. We refer to vulnerability features as the CVSS metrics of a single vulnerability $u_{i,i+1}\in U_{AP}$. Contrary, we refer to attack path features as the risk metrics for the attack path $AP$ (for the defined risk model they are likelihood and impact).
With this definition, we refer to attack path analysis as the answer to queries on attack path features. More formally:
@@
An attack path Query $Q$ is a query that specifies a range of values for one or more attack path features, potentially in conjunction or disjunction, to express a human analyst or an automatic process information need.
For example, the query $Q=\{impact~{}\geq~{}0.9\wedge likelihood<0.5\}$ represents the need to look at not probable ($<0.5$) but very dangerous ($\geq 0.9$) potential attacks.
Since generating an AG (i.e., enumerating all its paths) and analyzing attack paths is computationally expensive, we introduce a scalable framework in which the analysis can be performed before the completion of AG generation, and the generation can be steered by the analysis queries.
@@ 3 Overview of Our Approach
According to the classic process of attack path analysis, first defined by Phillips and Swiler [30] and currently used [8, 31, 32], it is composed of two principal milestones. Fig. 1-(a) depicts this classic workflow. We can recognize the first milestone that is the AG generation and the second one is the analysis of all its attack paths. The foundation of this framework is the representation of all the possible attack steps to provide a comprehensive view of the potential attacker’s movements on the network. Once the AG is generated, the subsequent step is the analysis of its paths, to study the attacks that can be put in place by an adversary. The problem of enumerating the attack paths is combinatorial in the number of edges, determining one of the main sources of their poor scalability. As a consequence, in the case of medium-large networks, the attack path analysis may never be performed because of this complexity.
@@
In this paper, we propose a new framework for AG generation and analysis that leverages the fundamental concepts of progressive data analysis [18, 19], that we report in Fig. 1-(b). The main idea behind progressive data analysis is to produce partial results during execution with increasing accuracy to generate intermediary outputs while the data is still being processed, potentially targeting interactive times. In the scenario of AG generation and analysis, the framework we propose considers the computation of early partial results that would give a coarse-grained approximation of the complete AG.
@@
At a certain instant $t_{significant}$ (driven by the amount of processed data), the partial AG becomes statistically significant (SPAG), in the sense that it is a good statistical approximation of the complete AG. Statistical significance must take into account the AG components that may vary in different networks or even that may change over time. For this reason, it must be evaluated on different attacker conditions (privileges on hosts) and vulnerability distribution. From this moment on, attack path analysis generates mature partial results, which reflect the final outcome within an acceptable margin of error, quantifiable through statistical indicators. This is the first advancement of the proposed framework, that can produce meaningful results to start exploring the attack paths at interactive time and continuously, allowing first observations and exploration without incurring the cost and waiting time of the full generation.
@@
As the progressive AG generation continues to run, there will be a certain instant $t_{stable}$ when the statistical significance reaches a threshold that guarantees definitive partial results. In these cases, the attack path analysis produces results that will no longer change substantially from the exact and, as such, supporting tasks like confirmatory analysis. This final part of the workflow runs until the complete AG is generated and/or attack paths are analyzed. At the same time, the approach hints at the possibility of leaving out this part, saving computational resources due to the diminishing return of the generation.
@@
This way of generating AGs allows the analysis of attack paths with preliminary data informed by quantifiable statistical indicators. However, given the exponential size of AGs with respect to the network and vulnerability inventory [9], it may still require a long time to reach the stability of definitive partial results. For this reason, the proposed framework contributes a steering mechanism to accelerate the convergence of the AG generation to the final outcome of any analysis executed during the generation [33]. This mechanism consists of informing the next step of generation with the current attack path analysis. For example, if the analysis query asks for the attack paths with the highest risk, then the steering mechanism prioritizes the generation of those paths that have high-risk values. It corresponds to the red blocks on the bottom of Fig. 1-(b).
@@
To automatically steer the generation with the analysis query we cannot directly query the AG generator, as it is not driven by attack path features. We leverage Machine Learning (ML) models to learn the vulnerability features from the structure of the attack paths and consequently prioritize the generation by selecting the vulnerability according to the learned features. For this reason, the first phase of AG generation must be agnostic to the attack path analysis (white blocks in Fig. 1-b) because it is used for collecting the initial balanced small set of attack paths that will be used to train the ML model and label attack paths as answers to the query or not. Consequently, the ML model is trained with the collected labeled attack paths to predict the vulnerability features that are representative only of attack paths that answer the query positively. Having obtained the intervals of vulnerability features that correspond to attack paths answering the query, they are used as input to the AG generator to prioritize the next generation of attack paths.
This steering mechanism allows a faster convergence of the partial AG to the portion of the complete AG that answers the attack path analysis query. In the next two sections, we describe the details of the statistically significant AG generation and the steering mechanism, along with the evaluation activities executed to demonstrate their validity.
4 StatAG: Statistically Significant Attack Graph Generation
@@
In this section, we describe how to achieve a progressively convergent statistically significant AG generation. It enables a new way of managing AGs with the main advantage of allowing the exploration and analysis of attack paths in partial AGs, without waiting for the complete generation, potentially very long due to its combinatorial complexity [8, 9]. The core idea of this approach, named Statistically Significant Attack Graph Generation (StatAG), is considering partial AGs being aware of statistical significance that becomes gradually higher until it reaches its maximum value corresponding to the complete AG (i.e., the exact result). Fig. 2 reports the workflow of StatAG.
Figure 2: Workflow for the statistically significant attack graph generation (StatAG).
The main activity during AG generation is the enumeration of its attack paths. While existing approaches [11, 34, 35] leverage common algorithms for graph traversal, i.e., BFS or DFS [36], we generate attack paths using random walks [37] over the reachability graph. A random walk of length $k$ is a stochastic process with random variables $X_{1},\cdots,X_{k}$ such that, given a graph $G=(V,E)$, $X_{i+1}$ is a vertex $v\in V$ chosen uniformly at random from the neighbors of $X_{i}$. The rationale for choosing random walks to visit the graph is due to their capabilities of generating unbiased samples [38], contrary to the classic traversal algorithms, which depend on the chosen starting node. As a consequence, random walks facilitate a quicker convergence of the path feature distributions of the partial AG to the complete AG [38] as we have experimentally proven in Appendix A.
@@
Following the sampling process, the next step involves constructing the corresponding attack path. Let us remember that the modeled AG is a multi-graph, therefore multiple edges (i.e., vulnerabilities) may exist between two nodes, although only one of the multi-edges can be part of an attack path, according to its definition (see Definition 2). When this happens, the construction of the attack path involves randomly selecting one vulnerability uniformly at random, so that the total number of attack paths aligns with the number of sampled walks.
The collection of attack paths constructed from the sampled walks defines what we name the partial AG.
@@
Let $G=(V,E)$ be a complete Attack Graph, then a Partial Attack Graph $PAG_{i}=(V_{i},E_{i})$ is an attack graph composed of a subset of nodes $V_{i}\subset V$ and edges $E_{i}\subset E$ retrieved after $i$ iterations of random walk sampling.
At this point, the approach iterates toward the generation of other partial AGs that are progressively added to reconstruct the complete AG, eventually. At the end of each iteration, we evaluate the statistical significance of the partial AG that quantifies the degree of its approximation to the complete AG. If it is acceptable, then the partial AG can be used for the initial exploration of the AG, along with indicators communicating its degree of uncertainty. To measure the statistical significance we need to define the null hypothesis $H_{0}$, which is the claim that no relationship exists between two sets of data being analyzed [39]. In the context of the proposed approach, the two sets of data are the attack path features of the partial AG and the complete AG, and the statistical significance refers to the probability of rejecting the null hypothesis. To evaluate this probability, we need to measure the probability $p$ of obtaining test results at least as extreme as the observed results (namely p-value) and the probability $\alpha$ of rejecting the null hypothesis (namely significance level [40]). According to these definitions, the partial AG is statistically significant when the $p\leq\alpha$ [41], with the value of $\alpha$ commonly set to 0.05 [39, 42].
We evaluate the null hypothesis through the Kolmogorov-Smirnov (KS) test [43], which is a statistical test to compare two data distributions and quantify the distance between them. In particular, we are interested in evaluating the distance between the attack path feature distributions of the complete and partial AGs. Let $\mathcal{D}_{AG}(x)$ and $\mathcal{D}_{PAG}(x)$ be the distribution of the attack path feature $x$ according to the complete and partial AG, respectively. Then, the KS distance between the two distributions is:
@@
where $sup_{x}$ is the supremum function [44], which corresponds to the least upper bound of the distances between the two distributions of $x$. Using the KS distance as defined in Equation 2, we can finally define the statistical significance of partial AGs.
@@
Let $AG$ be the complete attack graph and let $PAG_{i}$ be the partial attack graph generated at the $i^{th}$ iteration.
Let $H_{0}$: $KS(x)>T$ be the null hypothesis where $KS(x)$ is the KS distance between the distributions of the attack path feature $x$ for the complete and partial attack graph, and $T$ a predefined distance threshold.
Then, $PAG_{i}$ is statistically significant for the attack path feature $x$ if the $p\leq\alpha$, with $p$ the p-value of the $KS(x)$ and $\alpha$ the significance level.
Let us note that this definition is appropriate only for a posteriori evaluation of the statistical significance because it requires the complete attack graph $AG$. To enable the estimation of the statistical significance in a real-time application, we need a way to provide its indication during the progressive execution of the approach. For this purpose, we introduce the concept of Attack Path Feature stability (or simply stability) to quantify the variability of the partial results of an attack path feature gathered during the different iterations. To measure the stability of an attack path feature $x$, we consider the KS distance between the cumulative distribution of $x$ until the $i^{th}$ iteration and the one at iteration $i+1$. It indicates how much the new samples of iteration $i+1$ vary the already sampled distribution, with the rationale that a higher KS distance corresponds to more unstable results given the higher variability of the distribution.
@@
Let $PAG_{i}$ be the partial attack graph of the $i^{th}$ iteration of the approach and let $\mathcal{D}_{PAG_{i}}(x)$ be the cumulative distribution of the attack path feature $x$ in $PAG_{i}$. Then, the stability $\Delta$ of $x$ over $PAG_{i}$ is:

### v2

@@ Abstract
Attack Graph (AG) represents the best-suited solution to support cyber risk assessment for multi-step attacks on computer networks, although their generation suffers from poor scalability due to their combinatorial complexity. Current solutions propose to address the generation problem from the algorithmic perspective and postulate the analysis only after the generation is complete, thus implying too long waiting time before enabling analysis capabilities. Additionally, they poorly capture the dynamic changes in the networks due to long generation times. To mitigate these problems, this paper rethinks the classic AG analysis through a novel workflow in which the analyst can query the system anytime, thus enabling real-time analysis before the completion of the AG generation with quantifiable statistical significance. Further, we introduce a mechanism to accelerate the generation by steering it with the analysis query. To show the capabilities of the proposed framework, we perform an extensive quantitative validation and present a realistic case study on networks of unprecedented size. It demonstrates the advantages of our approach in terms of scalability and fitting to common attack path analyses.
Keywords:
Attack Graph Attack Path Analysis Progressive Computation Progressive Data Analysis Statistical Significance Computational Steering.
@@ 1 Introduction
In today’s digital world cyber attacks are becoming more complex to mitigate and have a higher impact on organizations’ infrastructure. In this scenario, it is crucial for any organization to promptly assess the potential cyber risks their networks are exposed to. This involves identifying the vulnerabilities in the network’s systems and quantifying the risk of their potential exploitation [9, 44]. Among the possible threat models, Attack Graph (AG) [17, 49] is a graph-based representation of the potential attack paths in a network used to analyze its cyber risk. It is currently the primary threat model used by enterprises for threat analysis and risk management [40]: among the others, Security Information and Event Management systems (e.g., IBM QRadar and LogRhythm) leverage AG for attack detection, network monitoring tools (e.g., Zeek and Suricata) integrate modules that generate AG to analyze malicious traffic, and threat intelligence platforms (e.g., ThreatConnect and OpenCTI) use AG to map out threat actor tactics, techniques, and procedures [27].
Although AGs are very expressive attack models, the above systems must face different problems. The first one is the poorly scalable computation of the attack paths, even for networks of moderate size, as AGs may grow exponentially in the size of network hosts and vulnerabilities [17]. Currently, much effort has been spent to find a trade-off between the accuracy of the model and computation time [48]. Indeed, providing scalable approaches for attack path analysis is still an open problem [33, 49]. A second emerging problem is that they postulate a classic analysis workflow in which the generation process comes first and the analysis comes next, aggravating the poor scalability of the generation process, as it slows down or completely stops the analysis in real scenarios [49]. A final problem concerns the alignment between the current network situation and the related AG model. While the classic workflow permits the attack path analysis on the complete set of attack steps, it also implies that any change in the network needs a re-computation of the whole attack graph to reflect such changes, exacerbating the scalability problem [9].
@@
To mitigate these problems, we contribute a novel workflow for AG generation and attack path analysis. It leverages the concept of statistical significance to express the degree of trustfulness of a partial AG, where not all the attack steps are computed. According to the foundational aspects of progressive data analysis [1, 5], it progressively supports attack path analysis already during the AG generation, allowing preventive analyses with quantitative quality-controlled statistically significant results. To improve the accuracy of analysis performed on partial AGs, we design a steering approach which automatically accelerates the AG generation based on the requested attack path analysis. Given a path analysis query from an analyst, it models the queried attack path features, extracts from them the features of each single attack step, and uses them to prioritize the generation of attack paths that answer the analysis query. We evaluate the speed-up of this process through a comprehensive scalability analysis performed on thousands of experiments, including different network scenarios and a real case study. Summarizing, this paper contributes: i) a novel framework for AG generation and attack path analysis based on progressive data analysis; ii) a way to express statistical significance over partial AGs; iii) a novel approach to steer the AG generation based on attack path analysis queries; iv) a comprehensive validation through experimental evaluations for both statistically significant AG generation and the steering mechanism; v) a case study demonstrating the possibility of managing large networks and conducting common attack path analyses. The source code and all materials are available as open source repository [^1] https://github.com/XAIber-lab/ProgressiveAttackGraph.
@@ 2 Preliminaries
Attack Graph Model. An Attack Graph (AG) is a graph-based representation of the possible paths an attacker can exploit to compromise an Information and Communications Technology (ICT) network. It is based on two inputs [14, 49]: the network reachability graph, and the network vulnerability inventory. The former is a directed graph where nodes are network hosts and edges are direct communication links between them. A vulnerability inventory reports the list of network hosts with an associated set of vulnerabilities, typically obtained by vulnerability scanners and security knowledge bases, as the National Vulnerability Database (NVD) [^2] https://nvd.nist.gov/. With these inputs, AG modeling represents all the dependencies between hosts and vulnerabilities an attacker can exploit. In this paper, we leverage the Topology Vulnerability Analysis (TVA) model [15], where the nodes are security conditions and represent the attacker access privileges in a specific host, while the edges are exploit dependencies and represent the possible movement of the attacker in case of a successful exploit.
@@
Once the AG model is defined, the next step is the Attack Graph Generation, which is the computation of all the attack paths to analyze the attack surface. They represent sequences of compromised devices and vulnerabilities exploited during the attack.
@@
Risk Model. When AG generation ends, cyber risks are associated with attack paths to analyze attack exposure or address specific requirements (e.g., damage to the network). They are evaluated according to existing approaches [9], that consider CVSS metrics [^3] https://www.first.org/cvss/v3.1/specification-document to estimate the likelihood and impact of an attack path $AP$ and calculate the risk according to its standard definition: $risk(AP)=likelihood(AP)\cdot impact(AP)$. The likelihood is calculated using the CVSS-3.1 exploitability metrics (Attack Vector, Attack Complexity, Privilege Required, and User Interaction). The impact is determined by CVSS-3.1 impact metrics. Likelihood, impact, and risk are in the range [0,1] (more details are available in the work by Gonzalez-Granadillo et al. [9]).
@@
Let $AP$ be an attack path and let $u_{1,2},\cdots\in U_{AP}$ the sequence of its vulnerabilities. We refer to vulnerability features as the CVSS metrics of a single vulnerability $u_{i,i+1}\in U_{AP}$. We refer to attack path features as the metrics for evaluating the risk of attack path $AP$ (for the defined risk model they are likelihood and impact).
Thus, an attack path analysis is the answer to queries on attack path features, e.g., $Q=\{impact~{}\geq~{}0.9\wedge likelihood<0.5\}$ represents the query for less probable ($<0.5$) but very dangerous ($\geq 0.9$) attacks. More formally:
@@
An attack path Query $Q$ is a query that specifies a range of values for one or more attack path features, potentially in conjunction or disjunction, to express an information need over AG.
@@ 3 Overview of Our Approach
Figure 1: (a) Classic attack graph generation and analysis process and (b) the progressive one. Dashed lines indicate periods in which the analysis is stalled.
@@
The classic process of attack path analysis [49], reported in Fig. 1-a, comprises two milestones: the AG generation and the analysis of its attack paths. The foundation of this framework is the representation of all the possible attack steps to provide a comprehensive view of the potential attacker’s movements on the network. The problem of enumerating the attack paths is combinatorial in the number of edges, determining one of the primary sources of their poor scalability. As a consequence, the attack path analysis may never be performed because of this complexity, even for medium-sized networks, or the analyst should make assumptions in the generation, limiting the expressiveness and validity of the results.
@@
In this paper, we propose a new framework for AG generation and analysis that leverages the fundamental concepts of progressive data analysis [1, 5], that we report in Fig. 1-b. The main idea is to produce partial results during execution with increasing accuracy to generate intermediary outputs, potentially targeting interactive times.
@@
In the case of AG generation and analysis, the framework we propose considers the computation of early partial results that would give a coarse-grained approximation of the complete AG. At a certain instant $t_{significant}$ (driven by the amount of processed data), the partial AG becomes statistically significant (StatAG), in the sense that it is a good statistical approximation of the complete AG. Statistical significance must consider the AG structure for its formulation. From that moment on, attack path analysis generates mature partial results, which reflect the outcome within an acceptable margin of error, quantifiable through statistical indicators. This is the first advancement of the proposed framework, producing meaningful results to continuously explore the attack paths at interactive time.
@@
As the progressive AG generation continues, there will be a certain instant $t_{stable}$ when the statistical significance reaches a threshold that guarantees definitive partial results. In these cases, the attack path analysis produces results that will no longer change substantially from the exact and, as such, supporting tasks like confirmatory analysis. This final part runs until the complete AG is generated or paths are analyzed.
@@
This way of generating AGs allows the analysis of attack paths with preliminary data informed by quantifiable statistical indicators. However, given the combinatorial size of AGs, it may still require a long time to reach the stability of definitive partial results. For this reason, we contribute a steering mechanism (SteerAG) to accelerate the convergence of the AG generation to the final outcome of any analysis executed during the generation [13]. It consists of informing the next step of generation with the current attack path analysis. For example, if the analysis query asks for the attack paths with the highest risk, then the steering mechanism prioritizes the generation of those high-risk paths.
@@
To automatically steer the generation with the analysis query we cannot directly query the AG generator, as attack path features do not drive it. We leverage Machine Learning (ML) models to learn the vulnerability features from the structure of the attack paths and consequently prioritize the generation by selecting the vulnerability according to the learned features. For this reason, the first phase of AG generation must be agnostic to the attack path analysis (white blocks in Fig. 1-b) because it is used for collecting the initial balanced small set of attack paths that will be used to train the ML model and label attack paths as answers to the query or not. Consequently, the ML model is trained with the collected labeled attack paths to predict the vulnerability features representative only of attack paths that answer the query positively. This steering mechanism allows a faster convergence of the partial AG to the portion of the complete AG that answers the attack path analysis query. In the next two sections, we describe the details of the statistically significant AG generation and the steering mechanism, along with their validation.
4 StatAG: Statistically Significant Generation
Figure 2: StatAG workflow.
@@ 4 StatAG: Statistically Significant Generation
This section describes how to achieve a progressively convergent statistically significant AG generation. It enables a new way of managing AGs with the main advantage of allowing the exploration and analysis of attack paths in partial AGs without waiting for the complete generation. The core idea of this approach, named Statistically Significant Attack Graph Generation (StatAG), is considering partial AGs being aware of statistical significance that becomes gradually higher until it reaches its maximum value corresponding to the complete AG (i.e., the exact result). Fig. 2 reports the workflow of StatAG. The main activity during AG generation is the enumeration of its attack paths. While existing approaches leverage common algorithms for graph traversal, i.e., BFS or DFS, we generate attack paths using random walks [25] over the reachability graph. A random walk is a path made of a succession of random steps in the graph. The rationale for random walks is their capability of generating unbiased samples, contrary to the classic traversal algorithms, which depend on the starting node. Thus, they facilitate a quicker convergence of the path features of the partial AG to the complete one [25].
@@
The next step involves constructing the corresponding attack path. Let us remember that the modeled AG is a multi-graph. Therefore, multiple edges (i.e., vulnerabilities) may exist between two nodes, although only one of the multi-edges can be part of an attack path, according to Definition 2. When this happens, the construction of the attack path involves randomly selecting one vulnerability uniformly at random so that the total number of attack paths aligns with the number of sampled walks. The collection of attack paths constructed from the sampled walks defines the partial AG.
@@
Let $G=(V,E)$ be a complete Attack Graph, then a Partial Attack Graph $PAG_{i}=(V_{i},E_{i})$ is composed of a subset of nodes $V_{i}\subset V$ and edges $E_{i}\subset E$ retrieved after $i$ iterations of random walk sampling.
At this point, the approach iterates toward the generation of other partial AGs that are progressively added to reconstruct the complete AG, eventually. At the end of each iteration, we evaluate the statistical significance of the partial AG that quantifies the degree of its approximation to the complete AG. If it is acceptable, then the partial AG can be used for the initial exploration of the AG, along with indicators communicating its degree of uncertainty. To measure the statistical significance, we need to define the null hypothesis $H_{0}$, which claims that no relationship exists between two sets of data being analyzed [39]. In the context of the proposed approach, the two sets of data are the attack path features of the partial AG and the complete AG, and the statistical significance refers to the probability of rejecting the null hypothesis. To evaluate this probability, we need to measure the probability $p$ of obtaining test results at least as extreme as the observed results (namely p-value) and the probability $\alpha$ of rejecting the null hypothesis (namely significance level). According to these definitions, the partial AG is statistically significant when the $p\leq\alpha$, with the value of $\alpha$ commonly set to 0.05 [39].
We evaluate the null hypothesis through the Kolmogorov-Smirnov (KS) statistical test, which compares two data distributions and quantifies the distance between them [30]. In particular, we evaluate the distance between the attack path feature distributions of the complete and partial AGs. Let $\mathcal{D}_{AG}(x)$ and $\mathcal{D}_{PAG}(x)$ be the distribution of the attack path feature $x$ according to the complete and partial AG, respectively. The KS distance of the two distributions is:
@@
where $sup_{x}$ is the supremum function, which corresponds to the least upper bound of the distances between the two distributions of $x$. According to this definition, we can define the statistical significance of partial AGs.
@@
Let $AG$ be the complete attack graph and let $PAG_{i}$ be the partial attack graph generated at the $i^{th}$ iteration. Let $H_{0}$: $KS(x)>T$ be the null hypothesis where $KS(x)$ is the KS distance between the distributions of the attack path feature $x$ for the complete and partial attack graph, and $T$ a predefined distance threshold. Then, $PAG_{i}$ is statistically significant for the attack path feature $x$ if the $p\leq\alpha$, with $p$ the p-value of the $KS(x)$ and $\alpha$ the significance level.
Let us note that this definition is appropriate only for a posteriori evaluation of the statistical significance because it requires the complete attack graph $AG$. To estimate the statistical significance in a real-time application, we need a way to provide its measure during the progressive execution of the approach. For this purpose, we introduce the concept of Attack Path Feature stability (or simply stability) to quantify the variability of the partial results of an attack path feature gathered during the different iterations. To measure the stability of an attack path feature $x$, we consider the KS distance between the cumulative distribution of $x$ until the $i^{th}$ iteration and the one at iteration $i+1$. It indicates how much the new samples of iteration $i+1$ vary the already sampled distribution, with the rationale that a higher KS distance corresponds to more unstable results given the higher variability.
@@
Let $PAG_{i}$ be the partial attack graph after $i$ iterations of the approach and $\mathcal{D}_{PAG_{i}}(x)$ the cumulative distribution of the attack path feature $x$ in $PAG_{i}$. The stability $\Delta$ of $x$ over $PAG_{i}$ is: